# Information Must Flow: Recursive Bootstrapping for Information Bottleneck in Optimal Transport

### v1

@@ Abstract
We introduce the Context-Content Uncertainty Principle (CCUP) as a unifying framework for understanding how information flows through cognitive systems by resolving the broken symmetry between high-entropy context and low-entropy content. Inference is framed as a dynamic cycle that aligns context with content through selective, bidirectional interactions, enabling information to propagate efficiently across representational hierarchies. Within this framework, cycle formation resolves the Information Bottleneck in Optimal Transport (iBOT) by coordinating bottom-up contextual disambiguation with top-down content reconstruction. This dual process is implemented via a Rao-Blackwellized inference strategy, formalized as variational entropy minimization, which steers the system toward minimal joint uncertainty while preserving the directionality of inference.
Crucially, local cycle completion gives rise to temporally extended chain formation (temporal bootstrapping) and spatially distributed hierarchical composition (spatial bootstrapping), enabling the system to simulate goals, evaluate counterfactuals, and iteratively refine internal models. To formalize this bootstrapped memory process, we prove a Delta Convergence Theorem showing that recursive entropy minimization over time yields structured memory states that converge to delta-like distributions over latent manifolds. These converged attractors act as fixed points in the flow of information, stabilizing perceptual schemas and motor plans as compact, generalizable priors. Spatial bootstrapping through ventral–dorsal functional segregation and temporal bootstrapping via the perception–action cycle support progressive refinement of sensorimotor models through continuous interaction with the environment. The sleep–wake cycle extends this dynamic across timescales, transforming episodic experiences into semantic knowledge via recursive consolidation, allowing the flow of information to bridge immediate perception with long-term abstraction.
@@ 1 Problem Formulation: What is iBOT and How to Tackle it?
Inverted inference gains its full expressive power when extended through recursive bootstrapping across both temporal and spatial hierarchies li2025Inverted . In the temporal domain, each content state is updated not in isolation but as part of a dynamic sequence, where structure learned at time $t$ serves as the scaffold for inference at $t+1$. This recursive alignment forms the basis of bootstrapped learning dynamics li2025CCUP , where the system iteratively refines its latent structure by minimizing conditional entropy over successive episodes. In parallel, spatial bootstrapping enables the composition of higher-level abstractions from structured subcomponents. Here, the structure is built recursively by assembling local regularities into increasingly abstract representations, satisfying both inference sufficiency and compositional coherence. The unification of these two bootstrapping processes, temporal and spatial, forms a multiscale inference engine wherein each layer constrains the next, and structure is not only inferred but grown, recursively and contextually, in space and time.
@@ 2 Theoretical Foundation: Context–Content Uncertainty Principle (CCUP)
[…]
@@
li2025Inverted Let $\Psi$ denote a high-entropy context variable and $\Phi$ a low-entropy content variable, with bidirectional inference maps $q(\Phi|\Psi)$ (inference) and $p(\Psi|\Phi)$ (generation). The pair $(\Psi,\Phi)$ is said to be aligned if the following equivalent conditions hold: 1) Cycle Consistency: $D_{\mathrm{KL}}(\Psi\,\|\,\hat{\Psi})\leq\varepsilon,D_{\mathrm{KL}}(\Phi\,\|\,\hat{\Phi})\leq\varepsilon$, where $\hat{\Phi}=q(\Phi|\Psi)$ and $\hat{\Psi}=p(\Psi|\hat{\Phi})$, and $\varepsilon\geq 0$ is a small alignment tolerance. 2) Mutual Information Saturation: $I(\Psi;\Phi)\approx H(\Phi)\approx H(\Psi)$, indicating that context and content share nearly all their informational degrees of freedom. 3) Entropy Symmetry (CCUP Condition): $H(\Phi|\Psi)\approx 0,H(\Psi|\Phi)\approx 0$, such that each variable fully determines the other under minimal uncertainty. 4) Optimal Transport Coupling: $\gamma^{*}=\arg\min_{\gamma\in\Pi(P_{\Psi},P_{\Phi})}\mathbb{E}_{\gamma}[c(\Psi,\Phi)]-\varepsilon H(\gamma)$, where $\gamma^{*}$ is a near-deterministic transport plan with low entropy, implying functional invertibility between context and content.
@@ 2.2 Dynamic Alignment Requires Inverted Inference and Bootstrapping
[…]
@@ 2.3 Bootstrapping as a Meta-Strategy for Learning and Inference
Let $\Phi\in\mathcal{S}$ denote structured content, and let $\Psi\in\mathcal{C}$ be contextual observations. Consider the latent variable $Z\in\mathcal{Z}$ mediating the generative path from structure to context: $p(Z|\Phi),\quad p(\Psi|Z)$. Define the forward generative prediction: $\hat{\Psi}=\mathbb{E}_{Z\sim p(Z|\Phi)}\left[\mathbb{E}_{\Psi\sim p(\Psi|Z)}[\Psi]\right]$, and define the backward recognition refinement via amortized inference: $\hat{Z}=q(Z|\Psi),\quad\Phi^{\prime}=q(\Phi|\hat{Z})$. Assume the update $\Phi\to\Phi^{\prime}$ minimizes a context-aligned prediction error: $\Phi^{\prime}=\arg\min_{\Phi}D(\Psi\,\|\,\hat{\Psi}(\Phi))$, where $D$ is a divergence metric (e.g., KL or squared error). Then the half-cycle operator: $\Phi\xrightarrow{p(Z|\Phi)\rightarrow p(\Psi|Z)}\hat{\Psi}\xrightarrow{q(Z|\hat{\Psi})\rightarrow q(\Phi^{\prime}|Z)}\Phi^{\prime}$ is equivalent to a predictive coding update that refines $\Phi$ by descending the gradient of prediction error on $\Psi$.
@@
Let $\Psi\in\mathbb{R}^{D}$ denote a high-dimensional context variable, and let $\Phi\in\mathbb{R}^{d}$ denote a structured latent representation with $d\ll D$. Define a latent bottleneck variable $Z\in\mathbb{R}^{k}$ with $d\leq k\ll D$, and consider a cycle-consistent inference scheme: $\Phi^{(t)}\xrightarrow{p(Z|\Phi)}Z^{(t)}\xrightarrow{p(\Psi|Z)}\hat{\Psi}^{(t)}\xrightarrow{q(Z|\hat{\Psi})}\hat{Z}^{(t)}\xrightarrow{q(\Phi|\hat{Z})}\Phi^{(t+1)}$. Assume 1) The joint model $p(\Psi,Z,\Phi)$ factorizes through low-dimensional conditional distributions; 2) The update $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\hat{\Psi}^{(t)})$ is entropy-decreasing: $H(\Psi|\Phi^{(t+1)})<H(\Psi|\Phi^{(t)})$; 3) The reconstruction steps $q(Z|\hat{\Psi})$, $q(\Phi|\hat{Z})$ are amortized and Lipschitz-continuous, then the inference process converges to a structured representation $\Phi^{\ast}$ that achieves predictive sufficiency: $H(\Psi|\Phi^{\ast})\approx\min_{\Phi}H(\Psi|\Phi)$ by avoiding direct marginalization over $\Psi\in\mathbb{R}^{D}$ and reducing inference complexity from $\mathcal{O}(\exp(D))$ to $\mathcal{O}(k^{2})$ under structured priors and amortization.
@@ 3 Delta Convergence and the Emergence of Generalizable Structure
Let $V_{t}\in\mathcal{X}$ be an evolving representation at iteration $t$, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a compact, convex space. Suppose the update rule takes the bootstrapped form: $V_{t+1}=\mathcal{F}(V_{t},\widetilde{E}(V_{t}))$, where $\mathcal{F}:\mathcal{X}\times\mathcal{Y}\to\mathcal{X}$ is a deterministic update operator, $\widetilde{E}:\mathcal{X}\to\mathcal{Y}$ is a surrogate or predictive operator and $\mathcal{F}$ is contractive in its first argument: there exists $10<\gamma<$ 1 such that $\|\mathcal{F}(x_{1},y)-\mathcal{F}(x_{2},y)\|\leq\gamma\|x_{1}-x_{2}\|\quad\text{for all }x_{1},x_{2}\in\mathcal{X},y\in\mathcal{Y}$. Then the sequence $\{V_{t}\}_{t=0}^{\infty}$ converges to a unique fixed point $V^{\ast}\in\mathcal{X}$, satisfying: $V^{\ast}=\mathcal{F}(V^{\ast},\widetilde{E}(V^{\ast}))$.
@@ 3.3 Variational Formulation as an Operational Principle of Inverted Inference
[…]
@@ 4.2 Hierarchical Composition as Spatial Bootstrapping
Let $\{\Psi_{\ell}\}_{\ell=0}^{L}$ be a hierarchy of latent variables forming a compositional structure under the conditions of Theorem LABEL:thm:HCSS. Suppose: 1) Each level $\Psi_{\ell}$ is obtained via entropy-aligned composition: $\Psi_{\ell}=\arg\min_{\Psi}\left[H(\Phi_{\ell-1}|\Psi)+\lambda_{\ell}D_{\mathrm{KL}}\left(\Psi\,\|\,f_{\ell}(\Psi_{\ell-1}^{(1)},\ldots,\Psi_{\ell-1}^{(n_{\ell})})\right)\right]$; 2) The conditional entropy decreases at each level: $H(\Phi_{\ell-1}|\Psi_{\ell})<H(\Phi_{\ell-1}|\Psi_{\ell-1}^{(i)}),\quad\forall i$; 3) The KL divergence across levels is uniformly bounded: $D_{\mathrm{KL}}\left(p(\Phi_{\ell-1}|\Psi_{\ell})\,\|\,p(\Phi_{\ell-1}|\Psi_{\ell-1}^{(i)})\right)\leq\epsilon_{\ell}$, then, in the limit $\ell\to L$, the top-level representation $\Psi_{L}$ converges in distribution to a delta-like measure concentrated around a minimal uncertainty manifold $\mathcal{M}^{\ast}\subset\mathcal{S}_{L}$, such that: $\lim_{\ell\to L}p(\Phi_{\ell-1}|\Psi_{\ell})\to\delta_{\mathcal{M}^{\ast}}(\Phi_{\ell-1})$ This implies that inference becomes maximally specific with respect to context, while remaining globally structured and compositionally abstract.
@@
[…]
@@ 5 Social Bootstrapping and the Origin of Language
From Spatiotemporal Bootstrapping to Social Bootstrapping.
@@
Let $\Phi\in\mathcal{S}$ denote the structured internal state (e.g., latent intent, belief, or concept) of a speaker, and let $\Psi\in\mathcal{C}$ be the observable signal (e.g., utterance, gesture, expression) transmitted during communication. Let $\hat{\Phi}\in\mathcal{S}$ denote the listener’s inferred internal state based on observing $\Psi$. The Information Bottleneck for Social Communication is the problem of finding a stochastic encoding $p(\Psi|\Phi)$ that minimizes the mutual information between the internal state and the signal, $\min_{p(\Psi|\Phi)}I(\Phi;\Psi)$, subject to a constraint on the relevance of the signal to the listener’s inference task: subject to $\text{subject to }I(\Psi;\hat{\Phi})\geq\mathcal{I}_{0}$, where $\mathcal{I}_{0}$ is a lower bound on the amount of relevant information that must be preserved for successful communication. This tradeoff captures the core challenge of social communication: compressing internal content into efficient signals that remain maximally informative under the listener’s reconstruction process.
@@
Let $\Phi\in\mathcal{S}$ be the structured internal content of a speaker (e.g., belief, intention), and $\Psi\in\mathcal{C}$ be the observable communicative signal (e.g., speech, gesture). Let $Z\in\mathcal{L}$ denote a shared latent linguistic representation. Language acts as an intermediating variable $Z$ that solves the Information Bottleneck (IB) for Optimal Transport (OT) in social communication, by satisfying the following: 1) Compression: The map $\Phi\to Z$ minimizes the encoding cost while preserving communicative relevance: $\min_{p(Z|\Phi)}I(\Phi;Z)\quad\text{subject to}\quad I(Z;\hat{\Phi})\geq\mathcal{I}_{0}$; 2) Transport: The signal $\Psi$ is generated from $Z$ in a way that minimizes semantic transport mismatch: $\min_{p(\Psi|Z)}D_{\mathrm{OT}}\big{(}p(\Psi|Z),\,p(\Psi|\hat{Z})\big{)}$; 3) Cycle-consistency: The listener infers $\hat{Z}\to\hat{\Phi}$, and successful communication occurs when $\hat{\Phi}\approx\Phi$, closing the inference loop. Hence, language defines a shared latent space $\mathcal{L}$ that compresses, transports, and reconstructs structured meaning across agents.
@@
$\Phi_{1}$$\Phi_{2}$$\Phi_{3}$hared Latent Language $Z\in\mathcal{L}$$\Psi_{1\to 2}$$\Psi_{2\to 3}$$\Psi_{3\to 1}$ulti-Agent Alignment via Language
@@
Let $\Phi=(\phi_{\text{manner}},\phi_{\text{path}},\phi_{\text{goal}})\in\mathcal{S}$ denote a structured latent meaning corresponding to a motion event, such as “The girl ran into the house.” This semantic tuple includes: 1) $\phi_{\text{manner}}$: the manner of motion (e.g., running); 2) $\phi_{\text{path}}$: the trajectory (e.g., entering); 3) $\phi_{\text{goal}}$: the spatial endpoint (e.g., the house). Let $Z=(z_{1},z_{2},z_{3})\in\mathcal{L}_{1}\times\mathcal{L}_{2}\times\mathcal{L}_{3}$ be a compositional symbolic representation of $\Phi$, where each $z_{i}$ is selected from a finite vocabulary slot corresponding to one semantic factor. Across languages, different mappings from $\Phi\to Z\to\Psi$ emerge, but each converges to a delta-like attractor within its symbolic system: 1) English (Satellite-framed): $z_{1}=\texttt{run}$, $z_{2}=\texttt{into}$, $z_{3}=\texttt{the house}$Rightarrow “The girl ran into the house.”

### v2

@@ Abstract
We introduce the Context-Content Uncertainty Principle (CCUP) as a unifying framework for understanding how information flows through cognitive systems by resolving the broken symmetry between high-entropy context and low-entropy content. Inference is framed as a dynamic cycle that aligns context with content through selective, bidirectional interactions, enabling information to propagate efficiently across representational hierarchies. Within this framework, cycle formation resolves the Information Bottleneck in Optimal Transport (iBOT) by coordinating bottom-up contextual disambiguation with top-down content reconstruction. This dual process is implemented via a memory-amortized inference (MAI) strategy, formalized as variational entropy minimization, which steers the system toward minimal joint uncertainty while preserving the directionality of inference.
Importantly, local cycle completion gives rise to temporally extended chain formation (temporal bootstrapping) and spatially distributed hierarchical composition (spatial bootstrapping), enabling the system to simulate goals, evaluate counterfactuals, and iteratively refine internal models. To formalize this bootstrapped memory process, we prove a Delta Convergence Theorem showing that recursive entropy minimization over time yields structured memory states that converge to delta-like distributions over latent manifolds. These converged attractors act as fixed points in the flow of information, stabilizing perceptual schemas and motor plans as compact, generalizable priors. Spatial bootstrapping through ventral–dorsal functional segregation and temporal bootstrapping via the perception–action cycle support progressive refinement of sensorimotor models through continuous interaction with the environment. The sleep–wake cycle extends this dynamic across timescales, transforming episodic experiences into semantic knowledge via recursive consolidation, allowing the flow of information to bridge immediate perception with long-term abstraction.
@@ 1 Problem Formulation: What is iBOT and How to Tackle it?
Inverted inference gains its full expressive power when extended through recursive bootstrapping across temporal and spatial hierarchies li2025Inverted . In the temporal domain, each content state is updated not in isolation but as part of a dynamic sequence, where structure learned at time $t$ serves as the scaffold for inference at $t+1$. This recursive alignment forms the basis of bootstrapped learning dynamics li2025CCUP , where the system iteratively refines its latent structure by minimizing conditional entropy over successive episodes. In parallel, spatial bootstrapping enables the composition of higher-level abstractions from structured subcomponents. Here, the structure is built recursively by assembling local regularities into increasingly abstract representations, satisfying both inference sufficiency and compositional coherence. The unification of these two bootstrapping processes forms a multiscale inference engine wherein each layer constrains the next, and structure is not only inferred but grown, recursively and contextually, in space and time.
@@ 2 Theoretical Foundation: Context–Content Uncertainty Principle (CCUP)
li2025CCUP Let $\Psi$ denote a high-entropy contextual variable and $\Phi$ denote a structured, low-entropy content variable such that their joint distribution $p(\Phi,\Psi)$ is well-defined. Then, for any inference system that seeks to minimize the uncertainty over $\Phi$ and $\Psi$, the optimal inference policy proceeds by: 1) Minimizing the joint entropy via decomposition: $H(\Phi,\Psi)=H(\Phi)+H(\Psi|\Phi)$; 2) Under the assumption $H(\Phi)\ll H(\Psi)$, the conditional entropy $H(\Psi|\Phi)$ dominates and becomes the principal target of minimization; 3) This induces a directional alignment: inference should proceed by first establishing $\Phi$, then using it to constrain $\Psi$, such that: $\Phi\leadsto\Psi\quad\text{(structure precedes specificity)}$. Moreover, the variational inference objective: $\mathcal{F}[q]=\mathbb{E}_{q(Z|\Psi)}[-\log p(\Psi|Z)]+D_{\mathrm{KL}}(q(Z|\Psi)\|p(Z|\Phi))$ is minimized when the prior $p(Z|\Phi)$ encodes low-entropy structure aligned with $\Psi$.
@@
li2025Inverted Let $\Psi$ denote a high-entropy context variable and $\Phi$ a low-entropy content variable, with bidirectional inference maps $q(\Phi|\Psi)$ (inference) and $p(\Psi|\Phi)$ (generation). The pair $(\Psi,\Phi)$ is said to be aligned if the following equivalent conditions hold: 1) Cycle Consistency: $D_{\mathrm{KL}}(\Psi\,\|\,\hat{\Psi})\leq\varepsilon,D_{\mathrm{KL}}(\Phi\,\|\,\hat{\Phi})\leq\varepsilon$, where $\hat{\Phi}=q(\Phi|\Psi)$ $\hat{\Psi}=p(\Psi|\hat{\Phi}), and$ $\varepsilon\geq 0$ is a small alignment tolerance. 2) Mutual Information Saturation: $I(\Psi;\Phi)\approx H(\Phi)\approx H(\Psi)$, indicating that context and content share nearly all their informational degrees of freedom. 3) Entropy Symmetry (CCUP Condition): $H(\Phi|\Psi)\approx 0,H(\Psi|\Phi)\approx 0$, such that each variable fully determines the other under minimal uncertainty. 4) Optimal Transport Coupling: $\gamma^{*}=\arg\min_{\gamma\in\Pi(P_{\Psi},P_{\Phi})}\mathbb{E}_{\gamma}[c(\Psi,\Phi)]-\varepsilon H(\gamma)$, where $\gamma^{*}$ is a near-deterministic transport plan with low entropy, implying functional invertibility between context and content.
@@ 2.2 Dynamic Alignment Requires Inverted Inference and Bootstrapping
To make inverted inference tractable and expressive, we introduce a latent variable $Z$ that serves as a variational mediator between structured content $\Phi$ and ambiguous context $\Psi$ (please refer to Fig. 1). Rather than modeling the direct mapping $p(\Psi|\Phi)$, which may be highly multi-modal or entangled, we define a two-stage generative model: $p(\Psi|\Phi)=\int p(\Psi|Z)\,p(Z|\Phi)\,dZ$, where $p(Z|\Phi)$ samples generalizable latent codes conditioned on compact structure and $p(\Psi|Z)$ decodes those codes into plausible contexts. The latent variable $Z$ functions as an information-theoretic stepping stool that facilitates the flow from abstract content $\Phi$ to entangled context $\Psi$. By mediating this transformation through an intermediate space, $Z$ enables a structured detour that preserves latent compositionality while overcoming the direct compression bottleneck imposed by high-dimensional context. This factorization decomposes the inference problem into two manageable steps: generalization via $p(Z|\Phi)$ and contextualization via $p(\Psi|Z)$, making efficient and precise structure-to-context alignment possible.
@@ 2.3 Bootstrapping as a Meta-Strategy for Learning and Inference
[…]
@@
[…]
@@ 3 Delta Convergence and the Emergence of Generalizable Structure
Let $V_{t}\in\mathcal{X}$ be an evolving representation at iteration $t$, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a compact, convex space. Suppose the update rule takes the bootstrapped form: $V_{t+1}=\mathcal{F}(V_{t},\widetilde{E}(V_{t}))$, where $\mathcal{F}:\mathcal{X}\times\mathcal{Y}\to\mathcal{X}$ is a deterministic update operator, $\widetilde{E}:\mathcal{X}\to\mathcal{Y}$ is a surrogate or predictive operator and $\mathcal{F}$ is contractive in its first argument: there exists $0<\gamma<1$ such that $\|\mathcal{F}(x_{1},y)-\mathcal{F}(x_{2},y)\|\leq\gamma\|x_{1}-x_{2}\|\quad\text{for all }x_{1},x_{2}\in\mathcal{X},y\in\mathcal{Y}$. Then the sequence $\{V_{t}\}_{t=0}^{\infty}$ converges to a unique fixed point $V^{\ast}\in\mathcal{X}$, satisfying: $V^{\ast}=\mathcal{F}(V^{\ast},\widetilde{E}(V^{\ast}))$.
@@ 3.3 Variational Formulation as an Operational Principle of Inverted Inference
Let $\Phi^{(t)}\in\mathcal{S}\subset\mathbb{R}^{d}$ denote the structured motor plan at time $t$, updated via recursive inference: $\Phi^{(t+1)}=\arg\min_{\Phi}\left[H(\Psi^{(t)}\mid\Phi)+\lambda\,\mathcal{D}(p(Z\mid\Phi),r(Z))\right]$, where $\Psi^{(t)}$ is the proprioceptive or external motor feedback, $Z$ is a latent motor intention, and $r(Z)$ is a prior over motor programs. In PD, due to dopaminergic degeneration and basal ganglia dysfunction, the recursive update operator becomes dysregulated and non-contractive: $\exists\,\Phi_{1},\Phi_{2}\in\mathcal{S},\quad\|\mathcal{F}(\Phi_{1},\hat{\Psi})-\mathcal{F}(\Phi_{2},\hat{\Psi})\|\geq\|\Phi_{1}-\Phi_{2}\|$, causing the sequence $\{\Phi^{(t)}\}$ to diverge or oscillate, leading to motor tremor and instability tarakad2019essential . Now, let a therapeutic intervention such as focused ultrasound (FUS) or deep brain stimulation (DBS) perlmutter2006deep modify the neural gain or effective damping of subcortical structures, particularly in the thalamus or subthalamic nucleus. This is modeled as restoring contractivity of the update operator $\mathcal{F}_{\mathrm{stim}}$: $\|\mathcal{F}_{\mathrm{stim}}(\Phi_{1},\hat{\Psi})-\mathcal{F}_{\mathrm{stim}}(\Phi_{2},\hat{\Psi})\|\leq\gamma\|\Phi_{1}-\Phi_{2}\|,\quad 0<\gamma<1$. Under these modified dynamics, the inference loop now satisfies the conditions for Delta Convergence: $\lim_{t\to\infty}p(Z\mid\Phi^{(t)})=\delta(Z-Z^{\ast})$, leading to stable convergence of the motor plan: $\Phi^{(t)}\to\Phi^{\ast},\quad\text{where }\Phi^{\ast}\text{ is a self-consistent, tremor-free attractor}$. From an inferential dynamics perspective, ultrasound and DBS therapies stabilize recursive updates in motor planning by restoring the contractive geometry of latent-to-context inference, allowing PD patients to recover precise, stable motor control.
@@ 4.2 Hierarchical Composition as Spatial Bootstrapping
Let $\{\Psi_{\ell}\}_{\ell=0}^{L}$ be a hierarchy of latent variables forming a compositional structure under the conditions of Theorem 3. Suppose: 1) Each level $\Psi_{\ell}$ is obtained via entropy-aligned composition: $\Psi_{\ell}=\arg\min_{\Psi}\left[H(\Phi_{\ell-1}|\Psi)+\lambda_{\ell}D_{\mathrm{KL}}\left(\Psi\,\|\,f_{\ell}(\Psi_{\ell-1}^{(1)},\ldots,\Psi_{\ell-1}^{(n_{\ell})})\right)\right]$; 2) The conditional entropy decreases at each level: $H(\Phi_{\ell-1}|\Psi_{\ell})<H(\Phi_{\ell-1}|\Psi_{\ell-1}^{(i)}),\quad\forall i$; 3) The KL divergence across levels is uniformly bounded: $D_{\mathrm{KL}}\left(p(\Phi_{\ell-1}|\Psi_{\ell})\,\|\,p(\Phi_{\ell-1}|\Psi_{\ell-1}^{(i)})\right)\leq\epsilon_{\ell}$, then, in the limit $\ell\to L$, the top-level representation $\Psi_{L}$ converges in distribution to a delta-like measure concentrated around a minimal uncertainty manifold $\mathcal{M}^{\ast}\subset\mathcal{S}_{L}$, such that: $\lim_{\ell\to L}p(\Phi_{\ell-1}|\Psi_{\ell})\to\delta_{\mathcal{M}^{\ast}}(\Phi_{\ell-1})$ This implies that inference becomes maximally specific with respect to context, while remaining globally structured and compositionally abstract.
@@
Let $\Phi\in\mathcal{S}\subset\mathbb{R}^{d}$ represent structured latent memory (e.g., semantic or episodic code), $Z\in\mathbb{R}^{k}$ an intermediate latent variable, and $\Psi\in\mathcal{C}\subset\mathbb{R}^{D}$ denote high-dimensional contextual observations. In the C2HM framework, memory is modeled as a recursive inference cycle: $\Phi^{(t)}\xrightarrow{p(Z|\Phi)}Z^{(t)}\xrightarrow{p(\Psi|Z)}\hat{\Psi}^{(t)}\rightarrow Z^{(t)\prime}\xrightarrow{q(\Phi|Z^{(t)\prime})}\Phi^{(t+1)}$, where the update $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\hat{\Psi}^{(t)})$ refines structured memory via simulation and correction. In Alzheimer’s disease (AD) scheltens2021alzheimer , this cycle is disrupted by multiple interacting mechanisms: 1) Accumulation of $\beta$-amyloid and tau impairs the structural integrity of $\mathcal{F}$, reducing plasticity and recursive refinement; 2) Network disconnection decouples the mapping $p(Z|\Phi)$, weakening simulation and pattern completion; 3) Chronic inflammation introduces stochastic noise in $p(\Psi|Z)$, increasing entropy in context predictions. Effective memory restoration, therefore, must target not just structural damage, but the recovery of the inference loop palop2016network . Let $\mathcal{T}=\{\text{Plasticity},\text{Stimulation},\text{Anti-Tau},\text{Simulation Training}\}$ denote a set of interventions. Each therapy $\tau\in\mathcal{T}$ acts to restore one or more components of the bootstrapping cycle. When these components are restored, the entropy of the forward prediction $H(\Psi|\Phi^{(t)})$ and the reconstruction divergence $D(\Phi^{(t+1)}\|\Phi^{(t)})$ both decrease over cycles. The system converges to a fixed point $\Phi^{\ast}$ such that: $\Phi^{\ast}=\arg\min_{\Phi}\mathbb{E}_{Z\sim p(Z|\Phi)}\left[-\log p(\Psi|Z)\right]+\lambda\cdot D_{\mathrm{KL}}(p(Z|\Phi)\|r(Z))$, where $r(Z)$ is a reference prior enforcing compactness. Therefore, memory restoration is best understood not as retrieval of stored data, but as recovery of an active, entropy-minimizing inference process.
@@ 5 Social Bootstrapping and the Origin of Language
5.1 From Spatiotemporal Bootstrapping to Social Bootstrapping
@@
Let $\Phi\in\mathcal{S}$ denote the structured internal state (e.g., latent intent, belief, or concept) of a speaker, and let $\Psi\in\mathcal{C}$ be the observable signal (e.g., utterance, gesture, expression) transmitted during communication. Let $\hat{\Phi}\in\mathcal{S}$ denote the listener’s inferred internal state based on observing $\Psi$. The Information Bottleneck for Social Communication is the problem of finding a stochastic encoding $p(\Psi|\Phi)$ that minimizes the mutual information between the internal state and the signal, $\min_{p(\Psi|\Phi)}I(\Phi;\Psi)$, subject to a constraint on the relevance of the signal to the listener’s inference task: $\text{subject to }I(\Psi;\hat{\Phi})\geq\mathcal{I}_{0}$, where $\mathcal{I}_{0}$ is a lower bound on the amount of relevant information that must be preserved for successful communication. This tradeoff captures the core challenge of social communication: compressing internal content into efficient signals that remain maximally informative under the listener’s reconstruction process.
@@
Let $\Phi\in\mathcal{S}$ be the structured internal content of a speaker (e.g., belief, intention), and $\Psi\in\mathcal{C}$ be the observable communicative signal (e.g., speech, gesture). Let $Z\in\mathcal{L}$ denote a shared latent linguistic representation. Language acts as an intermediating variable $Z$ that solves the Information Bottleneck (IB) for Optimal Transport (OT) in social communication, by satisfying the following: 1) Compression: The map $\Phi\to Z$ minimizes the encoding cost while preserving communicative relevance: $\min_{p(Z|\Phi)}I(\Phi;Z)\quad\text{subject to}\quad I(Z;\hat{\Phi})\geq\mathcal{I}_{0}$; 2) Transport: The signal $\Psi$ is generated from $Z$ in a way that minimizes semantic transport mismatch: $)\min_{p(\Psi|Z)}D_{\mathrm{OT}}\big{(}p(\Psi|Z),\,p(\Psi|\hat{Z}$$\big{)}$; 3) Cycle-consistency: The listener infers $\hat{Z}\to\hat{\Phi}$, and successful communication occurs when $\hat{\Phi}\approx\Phi$, closing the inference loop. Hence, language defines a shared latent space $\mathcal{L}$ that compresses, transports, and reconstructs structured meaning across agents.
@@
$\Phi_{1}$$\Phi_{2}$$\Phi_{3}$Shared Latent Language $Z\in\mathcal{L}$$\Psi_{1\to 2}$$\Psi_{2\to 3}$$\Psi_{3\to 1}$Multi-Agent Alignment via Language
@@
Let $\Phi=(\phi_{\text{manner}},\phi_{\text{path}},\phi_{\text{goal}})\in\mathcal{S}$ denote a structured latent meaning corresponding to a motion event, such as “The girl ran into the house.” This semantic tuple includes: 1) $\phi_{\text{manner}}$: the manner of motion (e.g., running); 2) $\phi_{\text{path}}$: the trajectory (e.g., entering); 3) $\phi_{\text{goal}}$: the spatial endpoint (e.g., the house). Let $Z=(z_{1},z_{2},z_{3})\in\mathcal{L}_{1}\times\mathcal{L}_{2}\times\mathcal{L}_{3}$ be a compositional symbolic representation of $\Phi$, where each $z_{i}$ is selected from a finite vocabulary slot corresponding to one semantic factor. Across languages, different mappings from $\Phi\to Z\to\Psi$ emerge, but each converges to a delta-like attractor within its symbolic system: 1) English (Satellite-framed): $z_{1}=\texttt{run}$, $z_{2}=\texttt{into}$, $z_{3}=\texttt{the house}$ $\Rightarrow$ “The girl ran into the house.”